# Estimates of automorphic forms on $\mathrm{SU}(n,1)$

### v1

@@ 3.2. Estimates of the Bergman metric
$\frac{1}{\big{(}1-|z|^{2}\big{)}^{a}}\bigg{|}\frac{1}{\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}\bigg{|}\cdot\bigg{|}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{i_{\alpha}}}\bigg{|}=\frac{\big{(}1-|z|^{2}\big{)}^{k(n+1)-a}}{\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{|}_{\mathrm{pet}}}\cdot\bigg{|}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{5/2}\big{)};$
$\frac{1}{\big{(}1-|z|^{2}\big{)}^{a}}\bigg{|}\frac{1}{\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}\bigg{|}\cdot\bigg{|}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial\overline{z}_{\beta}}\bigg{|}=\frac{\big{(}1-|z|^{2}\big{)}^{k(n+1)-a}}{\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{|}_{\mathrm{pet}}}\cdot\bigg{|}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial\overline{z}_{\beta}}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{a+5/2}\big{)},$
@@
(77) $\frac{1}{\big{(}1-|z|^{2}\big{)}^{a}}\bigg{|}\frac{1}{\big{(}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{)}^{2}}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial\overline{z}_{\beta}}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{a+5}\big{)}.$
Using estimates (3.10) and (77), for any $10\leq l\leq n-$ 1, we derive

### v2

@@ 3.2. Estimates of the Bergman metric
$\frac{1}{\big{(}1-|z|^{2}\big{)}^{a}}\bigg{|}\frac{1}{\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}\bigg{|}\cdot\bigg{|}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{i_{\alpha}}}\bigg{|}=$
$\frac{\big{(}1-|z|^{2}\big{)}^{k(n+1)-a}}{\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{|}_{\mathrm{pet}}}\cdot\bigg{|}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{n+a+2}\big{)};$
and
$\frac{1}{\big{(}1-|z|^{2}\big{)}^{a}}\bigg{|}\frac{1}{\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}\bigg{|}\cdot\bigg{|}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial\overline{z}_{\beta}}\bigg{|}=$
$\frac{\big{(}1-|z|^{2}\big{)}^{k(n+1)-a}}{\big{|}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{|}_{\mathrm{pet}}}\cdot\bigg{|}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial\overline{z}_{\beta}}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{n+a+2}\big{)},$
@@
(84) $\frac{1}{\big{(}1-|z|^{2}\big{)}^{a}}\bigg{|}\frac{1}{\big{(}\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)\big{)}^{2}}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial z_{\alpha}}\frac{\partial\operatorname{{\it{\mathcal{B}_{\Gamma}^{k(n+1)}}}}(z)}{\partial\overline{z}_{\beta}}\bigg{|}=O_{X_{\Gamma}}\big{(}k^{2n+a+4}\big{)}.$
Using estimates (3.10) and (84), for any $0\leq l\leq n-1$, we derive